OpenStudy (anonymous):

Find the product (8x^2-5y)(8x^2+5y)

OpenStudy (anonymous):

i think the answer would be 64x^4 - 25y^2

OpenStudy (anonymous):

yeah it is I have it broke down right here 64x^2-25y^2 = (8x)^2-(5y)^2 = (8x-5y)(8x+5y)

OpenStudy (anonymous):

approximate the area of the region R which is bounded above b the parbola y=x^2+1, below by the x-axis, on the left by the line x=-2 and on the right by the line x=3. can u find the answer?

OpenStudy (anonymous):

i think the best solution would be to integrate the function y, and then compute: Y(x=3)-Y(x=-2)

OpenStudy (anonymous):

64x^4-25y^2